# Thermodynamically consistent Cahn-Hilliard-Navier-Stokes equations using the metriplectic dynamics formalism

### 16 Feb 2024 v1

@@ 2.4.2 Fluxes and affinities for the GNS system
Comparison of (78) and (84) suggests we require the affinities associated with $\mathbf{m}$, $e$, and $\tilde{c}$. The conventional choices for these affinities are $\nabla(1/T)$, $\nabla(-\mathbf{v}/T)$, and $\nabla(-\mu/T)$, respectively (de Groot & Mazur, 1962). However, examination of (10) or (67) suggests using instead $\nabla T$, $\nabla\mathbf{v}$, and $\nabla\mu$, as was done in Coquinot & Morrison (2020).
@@
$L_{\mathbf{m}e}=T\bar{\bar{\Lambda}}\cdot\mathbf{v}\,,\quad L_{\mathbf{m}\mathbf{m}}=T\bar{\bar{\Lambda}}\,,\quad L_{\mathbf{m}\tilde{c}}=0\,,\quad L_{{e}\tilde{c}}=T\mu\bar{D}^{\prime}\,,$
@@
$S_{e}={1}/{T}\,,\quad S_{\mathbf{m}}=-{\mathbf{v}}/{T}\,,\quad\mathrm{and}\quad{S}_{\tilde{c}}=-{\mu}/{T}\,.$ (91)
Inserting these into $(o,S)_{H}$ using the 2-bracket of 89 yields the dissipative terms of (63), (65), and with the manipulations of transforming from $e$ to $\sigma$, those of the entropy equation (66). By direct calculation, as well as by construction, we obtain $(H,S)_{H}=0$ and $\dot{S}=({S},{S})_{H}=(S,H;S,H)\geq 0$ which reproduces (67).
To close the circle we transform the bracket of 89 in terms of the variables $(e,\mathbf{m},\tilde{c})$ to one in terms of $(\sigma,\mathbf{m},\tilde{c})$ via the following chain rule formulas:
$G_{e}\rightarrow G_{\sigma}/T,\quad G_{\mathbf{m}}\rightarrow G_{\mathbf{m}}-\mathbf{v}\,G_{\sigma}/{T},\quad\text{and}\quad G_{\tilde{c}}\rightarrow G_{\tilde{c}}-{\mu}\,G_{\sigma}/T\,.$ (92)
@@ 3.1 Hamiltonian and Entropy functional forms
The phenomenon of material transport along an interface is known as the Marangoni effect. The presence of a surface tension gradient naturally induces the migration of particles, moving from regions of low tension to those of high tension. This gradient can be triggered by a concentration gradient (or also a temperature gradient). In two-phase theory the interface between phases is regarded as being diffuse. According to the work of Taylor & Cahn (1998), one can model the diffuse interface by a single order parameter, say $\phi$ and with a free energy functional,
@@
In the isotropic surface energy case Guo & Lin (2015) develop a phase-field model for two-phase flow, which is thermodynamically consistent. The modeling based on a non-classical choice of energy and entropy, given respectively by
@@
where the coefficients $\epsilon_{S}$ and $\epsilon_{E}$ are assumed to be constant and $\Gamma$ is a homogeneous function of degree one that maps a vector to a real number.
@@
where $u$, $s$ and $f$ stand for the classical specific internal energy, entropy, and free energy, respectively, the coefficients $\lambda_{s}$ and $\lambda_{u}$ are constant parameters, and $\lambda_{f}(T)$ is a parameter depending on the temperature that will lead to anisotropic surface energy effects. He have defined the total densities $e^{a}_{\mathrm{Total}}$ and $\sigma^{a}_{\mathrm{Total}}$ for later use. The parameter $a$ takes on two values: $a=0$ reduces (98) and (99) to the expressions of Anderson et al. (2000), where we set $\epsilon^{2}_{E}=\lambda_{u}$ and $\epsilon^{2}_{S}=-\lambda_{s}$, while for $a=1$ they reduce to those used by Guo & Lin (2015) provided the choice of an isotropic surface energy is assumed, viz., $\Gamma(\nabla c)=|\nabla c|$. Thus, as is clear from (99), (98), and (100) that the dimensions of $\lambda_{f}$, $\lambda_{s}$, and $\lambda_{u}$ are either specific or volumetric depending on the case. As usual, we have the thermodynamic relation
@@
To summarize, our expressions (98) and (99) generalize the model studied by Guo & Lin (2015) by including $\Gamma$, which accounts for anisotropic surface energy effects, while our expressions generalize the model of Anderson et al. (2000) by including the factors of $\rho$ in the integrands making all quantities in the integrands specific quantities multiplied by the density, giving rise to more general sources of energy.
@@
$\mu^{a}_{\Gamma}\coloneqq\frac{\delta\mathfrak{F}^{a}}{\delta\tilde{c}}=\rho\frac{\partial u}{\partial\tilde{c}}-\nabla\cdot(\lambda_{f}\rho^{a}\Gamma\nabla c)=\mu-\frac{1}{\rho}\nabla\cdot(\lambda_{f}\rho^{a}\Gamma\bm{\xi})\,,$ (107)
@@
Upon setting $a=1$ (the case of Guo & Lin, 2015), this reduces to
@@
an experession that differs from that in Guo & Lin (2015) unless $\lambda_{f}\rho$ is constant. If this is the case and we choose a classical $\mu=c^{3}-c$, corresponding to the quartic Laudau potential, we obtain
@@
the chemical potential of Cahn and Hilliard who indeed make these assumptions (cf. page 267 of Cahn & Hilliard, 1958). For $\alpha=0$ (the case of Anderson et al., 2000), we have
@@ 3.2 Noncanonical Poisson bracket of the Cahn-Hilliard-Euler system
${\mathbf{m}}=\hat{\mathbf{m}}\,,\quad{\rho}=\hat{\rho}\,,\quad{\tilde{c}}=\hat{\tilde{c}}\,,\quad\sigma=\hat{\sigma}^{a}+\frac{\hat{\rho}^{a}}{2}\lambda_{s}\Gamma^{2}(\nabla\hat{c})\,,$ (112)
@@
$=\delta\sigma-\frac{1}{2}a\rho^{a-1}\lambda_{s}\Gamma^{2}(\nabla c)\delta\rho-\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\cdot\nabla\left(\frac{\delta\tilde{c}}{\rho}\right)+\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\cdot\nabla\left(\frac{\tilde{c}}{\rho^{2}}\delta\rho\right)\,,$ (115)
@@
$\int_{\Omega}{\hat{F}}_{\hat{\mathbf{m}}}\cdot\,\delta\hat{\mathbf{m}}+\hat{F}_{\hat{\rho}}\,{\delta\hat{\rho}}+\hat{F}_{\hat{\sigma}^{a}}\,\delta{\hat{\sigma}^{a}}+\hat{F}_{\hat{\tilde{c}}}\,\delta\hat{\tilde{c}}=\int_{\Omega}F_{\mathbf{m}}\cdot\delta\mathbf{m}+F_{\rho}\,\delta{\rho}+F_{\sigma}\,\delta{\sigma}+F_{\tilde{c}}\,\delta{\tilde{c}}\,.$ (116)
@@
$\{{F},G\}^{a}$ $=-\int_{\Omega}{\mathbf{m}}\cdot\left[F_{\mathbf{m}}\cdot\nabla G_{\mathbf{m}}-G_{\mathbf{m}}\cdot\nabla F_{\mathbf{m}}\right]$
$+{\rho}\left[F_{\mathbf{m}}\cdot\nabla\left(G_{\rho}-{a}\rho^{a-1}\lambda_{s}\Gamma^{2}G_{\sigma^{a}}/2-{\tilde{c}}\,\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}G_{\sigma^{a}}\right)/\rho^{2}\right)\right.$
$\hskip 28.45274pt-G_{\mathbf{m}}\cdot\nabla\left.\left(F_{\rho}-{a}\rho^{a-1}\lambda_{s}\Gamma^{2}F_{\sigma^{a}}/2-{\tilde{c}}\,\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}F_{\sigma^{a}}\right)/\rho^{2}\right)\right]$
$+\left({\sigma}^{a}+{{\rho}^{a}}\lambda_{s}\Gamma^{2}/2\right)\left[F_{\mathbf{m}}\cdot\nabla G_{\sigma^{a}}-G_{\mathbf{m}}\cdot\nabla F_{\sigma^{a}}\right]$
@@
$\hskip 28.45274pt-G_{\mathbf{m}}\cdot\nabla\left(F_{\tilde{c}}+\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}F_{\sigma^{a}}\right)/\rho\right)\big{]}\,.$ (118)
@@
$\frac{\delta S^{a}}{\delta\sigma^{a}}=1\,,\quad\frac{\delta S^{a}}{\delta\tilde{c}}=-\frac{1}{\rho}\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\right)\quad\frac{\delta S^{a}}{\delta\rho}=\frac{a}{2}\rho^{a-1}\lambda_{s}\Gamma^{2}+\frac{\tilde{c}}{\rho^{2}}\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\right)\,.$ (120)
@@
$H^{a}[\rho,\mathbf{m},\sigma,\tilde{c}]=\int_{\Omega}\frac{|\mathbf{m}|^{2}}{2\rho}+\rho u\left(\rho,\frac{\sigma^{a}}{\rho},\frac{\tilde{c}}{\rho}\right)+\frac{\rho^{a}}{2}\lambda_{u}\Gamma^{2}\left(\nabla\left(\frac{\tilde{c}}{\rho}\right)\right)\,.$ (122)
@@
$H^{a}_{\rho}$ $=-|\mathbf{v}|^{2}/2+u+{p}/{\rho}-sT-c\mu+a\rho^{a-1}\lambda_{u}\Gamma^{2}/2+{c}\,\nabla\cdot(\rho^{a}\lambda_{u}\Gamma{\bm{\xi}})/\rho\,,$
$H^{a}_{\tilde{c}}$ $=\mu-\nabla\cdot(\rho^{a}\lambda_{u}\Gamma{\bm{\xi}})\rho\,,$ (123)
@@
$\{F,G\}^{1}$ $=-\int_{\Omega}\mathbf{m}\,\cdot\left[F_{\mathbf{m}}\cdot\nabla G_{\mathbf{m}}-G_{\mathbf{m}}\cdot\nabla F_{\mathbf{m}}\right]$
$-\lambda_{s}\left[F_{\mathbf{m}}\cdot\nabla\cdot\left(\rho G_{\sigma^{1}}\Gamma{\bm{\xi}}\otimes\nabla c\right)-G_{\mathbf{m}}\cdot\nabla\cdot\left(\rho F_{\sigma^{1}}\Gamma{\bm{\xi}}\otimes\nabla c\right)\right]$ (124)
$+\sigma^{1}\left[F_{\mathbf{m}}\cdot\nabla G_{\sigma^{1}}-G_{\mathbf{m}}\cdot\nabla F_{\sigma^{1}}\right]+\tilde{c}\left[F_{\mathbf{m}}\cdot\nabla G_{\tilde{c}}-G_{\mathbf{m}}\cdot\nabla F_{\tilde{c}}\right]\,,$
@@
$\partial_{t}\mathbf{v}$ =$=\{\mathbf{v},H^{1}\}^{1}=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot(p\bar{I}+\rho\lambda_{f}\Gamma{\bm{\xi}}\otimes\nabla c)\,,$ (125)
@@
$\partial_{t}\sigma^{1}_{\mathrm{Total}}$ =$=\{{\sigma}^{1}_{\mathrm{Total}},H^{1}\}^{1}=-\mathbf{v}\cdot\nabla{\sigma}^{1}_{\mathrm{Total}}-{\sigma}^{1}_{\mathrm{Total}}\,\nabla\cdot\mathbf{v}\,,$ (128)
@@
$\{F,G\}^{0}$ $=-\int_{\Omega}\mathbf{m}\cdot\,\left[F_{\mathbf{m}}\cdot\nabla G_{\mathbf{m}}-G_{\mathbf{m}}\cdot\nabla F_{\mathbf{m}}\right]$
$-\lambda_{s}\left[F_{\mathbf{m}}\cdot\nabla\cdot\left(G_{\sigma^{0}}\Gamma{\bm{\xi}}\otimes\nabla c\right)-G_{\mathbf{m}}\cdot\nabla\cdot\left(F_{\sigma^{0}}\Gamma{\bm{\xi}}\otimes\nabla c\right)\right]$
@@
$+\sigma^{0}\left[F_{\mathbf{m}}\cdot\nabla G_{\sigma^{0}}-G_{\mathbf{m}}\cdot\nabla F_{\sigma^{0}}\right]+\tilde{c}\left[F_{\mathbf{m}}\cdot\nabla G_{\tilde{c}}-G_{\mathbf{m}}\cdot\nabla F_{\tilde{c}}\right]\,.$ (129)
@@
$\partial_{t}\mathbf{v}$ =$=\{\mathbf{v},H^{0}\}^{0}=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot\left[\left(p-\lambda_{f}\Gamma^{2}/2\right)\bar{I}+\lambda_{f}\Gamma{\bm{\xi}}\otimes\nabla c\right]\,,$ (130)
@@
$\partial_{t}\sigma^{0}_{\mathrm{Total}}$ =$=\{{\sigma}^{0}_{\mathrm{Total}},H^{0}\}^{0}=-\mathbf{v}\cdot\nabla{\sigma}_{\mathrm{Total}}^{0}-{\sigma}_{\mathrm{Total}}^{0}\,\nabla\cdot\mathbf{v}\,.$ (133)
Let us now comment on these two Hamiltonian systems. By construction both the $a=1$ and $a=0$ systems conserve their Hamiltonians and entropies, as given by (99) and (98) with $a=1$ and $a=0$ , respectively. Both systems have momentum equations containing a term describing anisotropic surface energy (capillary) effects. The $a=0$ system of (130)–(133) is identical to the ideal limit of that given in the work of Anderson et al. (2000). Upon choosing $\Gamma(\nabla c)=|\nabla c|$, the $a=1$ system of (125)–(128) should correspond to the ideal limit of that of Guo & Lin (2015), but it does not. In fact the system of Guo & Lin (2015) in this limit does not conserve energy. Moreover, the capillary effect in their momentum equation (equation (3.40)), which should be replaced by (125) with $\Gamma(\nabla c)=|\nabla c|$, vanishes in the one-dimensional limit. Since such surface effects are determined by mean or weighted mean curvature (Taylor, 1992), it is clear that this is physically untenable. Fortunately, our method provides a simple fix to their equations, while showing how to generalize them to include anisotropic surface effects.

### 29 Feb 2024 v2

@@ 2.4.2 Fluxes and affinities for the GNS system
Comparison of (77) and (83) suggests we require the affinities associated with $\mathbf{m}$, $e$, and $\tilde{c}$. The conventional choices for these affinities are $\nabla(1/T)$, $\nabla(-\mathbf{v}/T)$, and $\nabla(-\mu/T)$, respectively (de Groot and Mazur, 1962). However, examination of (10) or (66) suggests using instead $\nabla T$, $\nabla\mathbf{v}$, and $\nabla\mu$, as was done in Coquinot and Morrison (2020).
@@
$L_{\mathbf{m}e}=T\bar{\bar{\Lambda}}\cdot\mathbf{v}\,,\ \ \ L_{\mathbf{m}\mathbf{m}}=T\bar{\bar{\Lambda}}\,,\ \ \ L_{\mathbf{m}\tilde{c}}=0\,,\ \ \ L_{{e}\tilde{c}}=T\mu\bar{D}^{\prime}\,,$
@@
$S_{e}={1}/{T}\,,\ \ S_{\mathbf{m}}=-{\mathbf{v}}/{T}\,,\ \ \mathrm{and}\ \ {S}_{\tilde{c}}=-{\mu}/{T}\,.$ (90)
Inserting these into $(o,S)_{H}$ using the 2-bracket of 88 yields the dissipative terms of (62), (64), and with the manipulations of transforming from $e$ to $\sigma$, those of the entropy equation (65). By direct calculation, as well as by construction, we obtain $(H,S)_{H}=0$ and $\dot{S}=({S},{S})_{H}=(S,H;S,H)\geq 0$ which reproduces (66).
To close the circle we transform the bracket of 88 in terms of the variables $(e,\mathbf{m},\tilde{c})$ to one in terms of $(\sigma,\mathbf{m},\tilde{c})$ via the following chain rule formulas:
$G_{e}$ $\rightarrow G_{\sigma}/T,\qquad G_{\mathbf{m}}\rightarrow G_{\mathbf{m}}-\mathbf{v}\,,G_{\sigma}/{T}\,,$
$\quad G_{\tilde{c}}\rightarrow G_{\tilde{c}}-{\mu}\,G_{\sigma}/T\,.$ (91)
@@ 3.1 Hamiltonian and Entropy functional forms
The phenomenon of material transport along an interface is known as the Marangoni effect. The presence of a surface tension gradient naturally induces the migration of particles, moving from regions of low tension to those of high tension. This gradient can be triggered by a concentration gradient (or also a temperature gradient). In two-phase theory the interface between phases is regarded as being diffuse. According to the work of Taylor and Cahn (1998), one can model the diffuse interface by a single order parameter, say $\phi$ and with a free energy functional,
@@
In the isotropic surface energy case Guo and Lin (2015) develop a phase-field model for two-phase flow, which is thermodynamically consistent. The modeling based on a non-classical choice of energy and entropy, given respectively by
@@
where the coefficients $\epsilon_{S}$ and $\epsilon_{E}$ are assumed to be constant and $\Gamma$ is a homogeneous function of degree one that takes a vector to a real number.
@@
where $u$, $s$ and $f$ stand for the classical specific internal energy, entropy, and free energy, respectively, the coefficients $\lambda_{s}$ and $\lambda_{u}$ are constant parameters, and $\lambda_{f}(T)$ is a parameter depending on the temperature that will lead to anisotropic surface energy effects. He have defined the total densities $e^{a}_{\mathrm{Total}}$ and $\sigma^{a}_{\mathrm{Total}}$ for later use. The parameter $a$ takes on two values: $a=0$ reduces (97) and (98) to the expressions of Anderson et al. (2000), where we set $\epsilon^{2}_{E}=\lambda_{u}$ and $\epsilon^{2}_{S}=-\lambda_{s}$, while for $a=1$ they reduce to those used by Guo and Lin (2015) provided the choice of an isotropic surface energy is assumed, viz., $\Gamma(\nabla c)=|\nabla c|$. Thus, as is clear from (98), (97), and (99) that the dimensions of $\lambda_{f}$, $\lambda_{s}$, and $\lambda_{u}$ are either specific or volumetric depending on the case. As usual, we have the thermodynamic relation
@@
To summarize, our expressions (97) and (98) generalize the model studied by Guo and Lin (2015) by including $\Gamma$, which accounts for anisotropic surface energy effects, while our expressions generalize the model of Anderson et al. (2000) by including the factors of $\rho$ in the integrands making all quantities in the integrands specific quantities multiplied by the density, giving rise to more general sources of energy.
@@
$\mu^{a}_{\Gamma}$ $\coloneqq\frac{\delta\mathfrak{F}^{a}}{\delta\tilde{c}}=\rho\frac{\partial u}{\partial\tilde{c}}-\nabla\cdot(\lambda_{f}\rho^{a}\Gamma\nabla c)$
$=\mu-\frac{1}{\rho}\nabla\cdot(\lambda_{f}\rho^{a}\Gamma\bm{\xi})\,,$ (106)
@@
Upon setting $a=1$ (the case of Guo and Lin, 2015), this reduces to
@@
an expression that differs from that in Guo and Lin (2015) unless $\lambda_{f}\rho$ is constant. If this is the case and we choose a classical $\mu=c^{3}-c$, corresponding to the quartic Laudau potential, we obtain
@@
the chemical potential of Cahn and Hilliard who indeed make these assumptions (cf. page 267 of Cahn and Hilliard, 1958). For $\alpha=0$ (the case of Anderson et al., 2000), we have
@@ 3.2 Noncanonical Poisson bracket of the Cahn-Hilliard-Euler system
${\mathbf{m}}$ $=\hat{\mathbf{m}}\,,\quad{\rho}=\hat{\rho}\,,\quad{\tilde{c}}=\hat{\tilde{c}}\,,$
$\sigma$ $=\hat{\sigma}^{a}+\frac{\hat{\rho}^{a}}{2}\lambda_{s}\Gamma^{2}(\nabla\hat{c})\,,$ (111)
@@
$=\delta\sigma-\frac{1}{2}a\rho^{a-1}\lambda_{s}\Gamma^{2}(\nabla c)\delta\rho-\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\cdot\nabla\left(\frac{\delta\tilde{c}}{\rho}\right)$
$\hskip 56.9055pt+\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\cdot\nabla\left(\frac{\tilde{c}}{\rho^{2}}\delta\rho\right)\,,$ (114)
@@
$\int_{\Omega}{\hat{F}}_{\hat{\mathbf{m}}}\cdot\,\delta\hat{\mathbf{m}}$ $+\hat{F}_{\hat{\rho}}\,{\delta\hat{\rho}}+\hat{F}_{\hat{\sigma}^{a}}\,\delta{\hat{\sigma}^{a}}+\hat{F}_{\hat{\tilde{c}}}\,\delta\hat{\tilde{c}}$ (115)
$=\int_{\Omega}F_{\mathbf{m}}\cdot\delta\mathbf{m}+F_{\rho}\,\delta{\rho}+F_{\sigma}\,\delta{\sigma}+F_{\tilde{c}}\,\delta{\tilde{c}}\,.$
@@
$\{{F},G\}^{a}$ $=-\int_{\Omega}{\mathbf{m}}\cdot\Big{[}F_{\mathbf{m}}\cdot\nabla G_{\mathbf{m}}-G_{\mathbf{m}}\cdot\nabla F_{\mathbf{m}}\Big{]}$
$+{\rho}\,\Big{[}F_{\mathbf{m}}\cdot\nabla\big{(}G_{\rho}-{a}\rho^{a-1}\lambda_{s}\Gamma^{2}G_{\sigma^{a}}/2$
$\hskip 56.9055pt-{\tilde{c}}\,\nabla\cdot(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}G_{\sigma^{a}})/\rho^{2}\big{)}$
$\hskip 14.22636pt-G_{\mathbf{m}}\cdot\nabla\big{(}F_{\rho}-{a}\rho^{a-1}\lambda_{s}\Gamma^{2}F_{\sigma^{a}}/2$ (117)
$\hskip 56.9055pt-{\tilde{c}}\,\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}F_{\sigma^{a}}\right)/\rho^{2}\big{)}\Big{]}$
$+\left({\sigma}^{a}+{{\rho}^{a}}\lambda_{s}\Gamma^{2}/2\right)\Big{[}F_{\mathbf{m}}\cdot\nabla G_{\sigma^{a}}-G_{\mathbf{m}}\cdot\nabla F_{\sigma^{a}}\Big{]}$
@@
$\hskip 14.22636pt-G_{\mathbf{m}}\cdot\nabla\left(F_{\tilde{c}}+\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}F_{\sigma^{a}}\right)/\rho\right)\Big{]}\,.$
@@
$\frac{\delta S^{a}}{\delta\sigma^{a}}$ $=1\,,\quad\frac{\delta S^{a}}{\delta\tilde{c}}=-\frac{1}{\rho}\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\right)$
$\frac{\delta S^{a}}{\delta\rho}$ $=\frac{a}{2}\rho^{a-1}\lambda_{s}\Gamma^{2}+\frac{\tilde{c}}{\rho^{2}}\nabla\cdot\left(\rho^{a}\lambda_{s}\Gamma{\bm{\xi}}\right)\,.$ (119)
@@
$H^{a}[\rho,\mathbf{m},\sigma,\tilde{c}]$ $=\int_{\Omega}\frac{|\mathbf{m}|^{2}}{2\rho}+\rho u\left(\rho,\frac{\sigma^{a}}{\rho},\frac{\tilde{c}}{\rho}\right)$
$\hskip 45.52458pt+\frac{\rho^{a}}{2}\lambda_{u}\Gamma^{2}\left(\nabla\left(\frac{\tilde{c}}{\rho}\right)\right)\,.$ (121)
@@
$H^{a}_{\rho}$ $=-|\mathbf{v}|^{2}/2+u+{p}/{\rho}-sT-c\mu$
$\hskip 28.45274pt+a\rho^{a-1}\lambda_{u}\Gamma^{2}/2+{c}\,\nabla\cdot(\rho^{a}\lambda_{u}\Gamma{\bm{\xi}})/\rho\,,$
$H^{a}_{\tilde{c}}$ $=\mu-\nabla\cdot(\rho^{a}\lambda_{u}\Gamma{\bm{\xi}})/\rho\,,$ (122)
@@
$\{F,G\}^{1}$ $=-\int_{\Omega}\mathbf{m}\,\cdot\big{[}F_{\mathbf{m}}\cdot\nabla G_{\mathbf{m}}-G_{\mathbf{m}}\cdot\nabla F_{\mathbf{m}}\big{]}$
$+\rho\big{[}F_{\mathbf{m}}\cdot\nabla G_{\rho}-G_{\mathbf{m}}\cdot\nabla F_{\rho}\big{]}$
$-\lambda_{s}\big{[}F_{\mathbf{m}}\cdot\nabla\cdot\left(\rho G_{\sigma^{1}}\Gamma{\bm{\xi}}\otimes\nabla c\right)$
$\hskip 56.9055pt-G_{\mathbf{m}}\cdot\nabla\cdot\left(\rho F_{\sigma^{1}}\Gamma{\bm{\xi}}\otimes\nabla c\right)\big{]}$
$+\sigma^{1}\left[F_{\mathbf{m}}\cdot\nabla G_{\sigma^{1}}-G_{\mathbf{m}}\cdot\nabla F_{\sigma^{1}}\right]$
$+\tilde{c}\left[F_{\mathbf{m}}\cdot\nabla G_{\tilde{c}}-G_{\mathbf{m}}\cdot\nabla F_{\tilde{c}}\right]\,,$ (123)
@@
$\partial_{t}\mathbf{v}$ =$=\{\mathbf{v},H^{1}\}^{1}$
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot(p\bar{I}+\rho\lambda_{f}\Gamma{\bm{\xi}}\otimes\nabla c)\,,$ (124)
@@
$\partial_{t}\sigma^{1}_{\mathrm{Total}}$ =$=\{{\sigma}^{1}_{\mathrm{Total}},H^{1}\}^{1}$
$=-\mathbf{v}\cdot\nabla{\sigma}^{1}_{\mathrm{Total}}-{\sigma}^{1}_{\mathrm{Total}}\,\nabla\cdot\mathbf{v}\,,$ (127)
@@
$\{F,G\}^{0}$ $=-\int_{\Omega}\mathbf{m}\cdot\,\big{[}F_{\mathbf{m}}\cdot\nabla G_{\mathbf{m}}-G_{\mathbf{m}}\cdot\nabla F_{\mathbf{m}}\big{]}$
$+\rho\big{[}F_{\mathbf{m}}\cdot\nabla G_{\rho}-G_{\mathbf{m}}\cdot\nabla F_{\rho}\big{]}$
$-\lambda_{s}\big{[}F_{\mathbf{m}}\cdot\nabla\cdot\left(G_{\sigma^{0}}\Gamma{\bm{\xi}}\otimes\nabla c\right)$
$\hskip 68.28644pt-G_{\mathbf{m}}\cdot\nabla\cdot\left(F_{\sigma^{0}}\Gamma{\bm{\xi}}\otimes\nabla c\right)\big{]}$
@@
$+\sigma^{0}\left[F_{\mathbf{m}}\cdot\nabla G_{\sigma^{0}}-G_{\mathbf{m}}\cdot\nabla F_{\sigma^{0}}\right]$
$+\tilde{c}\left[F_{\mathbf{m}}\cdot\nabla G_{\tilde{c}}-G_{\mathbf{m}}\cdot\nabla F_{\tilde{c}}\right]\,.$ (128)
@@
$\partial_{t}\mathbf{v}$ =$=\{\mathbf{v},H^{0}\}^{0}$ (129)
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\frac{1}{\rho}\nabla\cdot\Big{[}\left(p-\lambda_{f}\Gamma^{2}/2\right)\bar{I}$
$\hskip 92.47145pt+\lambda_{f}\Gamma{\bm{\xi}}\otimes\nabla c\Big{]}\,,$
@@
$\partial_{t}\sigma^{0}_{\mathrm{Total}}$ =$=\{{\sigma}^{0}_{\mathrm{Total}},H^{0}\}^{0}$
$=-\mathbf{v}\cdot\nabla{\sigma}_{\mathrm{Total}}^{0}-{\sigma}_{\mathrm{Total}}^{0}\,\nabla\cdot\mathbf{v}\,.$ (132)
Let us now comment on these two Hamiltonian systems. By construction both the $a=1$ and $a=0$ systems conserve their Hamiltonians and entropies, as given by (98) and (97) with $a=1$ and $a=0$ , respectively. Both systems have momentum equations containing a term describing anisotropic surface energy (capillary) effects. The $a=0$ system of (129)–(132) is identical to the ideal limit of that given in the work of Anderson et al. (2000). Upon choosing $\Gamma(\nabla c)=|\nabla c|$, the $a=1$ system of (124)–(127) should correspond to the ideal limit of that of Guo and Lin (2015), but it does not. In fact the system of Guo and Lin (2015) in this limit does not conserve energy. Moreover, the capillary effect in their momentum equation (equation (3.40)), which should be replaced by (124) with $\Gamma(\nabla c)=|\nabla c|$, vanishes in the one-dimensional limit. Since such surface effects are determined by mean or weighted mean curvature (Taylor, 1992), it is clear that this is physically untenable. Fortunately, our method provides a simple fix to their equations, while showing how to generalize them to include anisotropic surface effects.